Evelyn bought stock in a company two years ago that was worth x dollars. During the first year that she owned the stock, it decreased by 36%. During the second year the value of the stock decreased by 20% Write an expression in terms of that represents the value of the stock after the two years have passed.

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52787)

.

The expression for the cost after 2 years is


    cost = x*(1-0.36)*(1-0.20) = x*0.64*0.8 = 0.512x.    ANSWER

Solved, answered, explained, calculated and completed.



Answer by greenestamps(13200)